# Perceptually Optimized Super Resolution

### v1

@@ 6.2 Subjective Quality Results
Figure 7: Our method quality map prediction results for ×$\times 4$ upsampling.
@@
7 Extension - AR/VR Display
Contrast sensitivity models such as the StelaCSF appropriately model human contrast perception as a function of eccentricity, and thus can extend our model to account for acuity across the visual field. Modern VR/AR headsets have built in eye-trackers that can be used to control our framework. In Fig. 9, we present some preliminary results for our quality map estimation with different gaze positions on the screen. The top row shows the eccentricity map relative to the gaze location, and the bottom row shows how our prediction for required SR quality varies. As anticipated, our perceptual model predicts that higher quality resolution will be used when the user is looking, while for areas in the periphery, our model predicts that the lowest up-sampling quality will be used. The main application is rendering in a lower resolution throughout the field of view, and then up-sampling the rendering for real-time VR/AR displays using our technique. This would especially be beneficial for high quality real-time streaming on AR/VR displays, where the gaze position is not known during transmission and traditional foveated rendering systems would fail, and transmitting a uniformly lower resolution steam and subsequent gaze-contingent up-sampling would be ideal.
Figure 9: Our model predictions based on gaze position with ×$\times 4$ super-resolution. In the first column, we have the original image and the corresponding quality map. In the other columns we have on top the eccentricity map expressed in degrees and bottom we have the corresponding quality map.
8 Future work
Although the method demonstrated expected performance with regard to video content, it was observed that aliasing issues were present. This phenomenon is frequently observed when a single frame is considered in isolation. It would be of interest to investigate the potential of our perceptual method in conjunction with SR techniques that have been developed explicitly for use in video applications.
The method can be extended to handle color perception be employing the VDP modeling such as ColorVDP [29]. The could enable applications to problems such was deep learning for color constancy/automatic-white-balancing.
Though we applied our perceptual model to control CNN based SR, a similar strategy can be employed for CNN based denoising, frame-interpolation and other low level image/video enhancement tasks. Some modifications to the perceptual modeling may need to be done in a task specific manner. For example, some assumptions about the spectral nature of the noise in denoising, or some pre-processing.
The most interesting open question and future direction is the perceptual control of modern generative models to make them efficient without noticeable loss in visual quality. It is likely that reducing computational complexity in techniques such as latent diffusion would result in similar fourier characteristics of the output i.e a more efficient pipeline (e.g quantized, depth reduction, less channels etc) would result in loss of higher frequencies and our model and pipeline can be applied by studying the attenuation curves similarly. The unclear part is that it might also lead to some structural degradation i.e shape distortion, which may would require some additional perceptual modeling as most perceptual models like the VDP rely on pixel wise comparisons. It would be an exciting problem of how the model can be extended to cater for the human perception of such degradations.
9 Conclusions
Contemporary work on image enhancement evaluates the quality of the results in terms of metrics such as PSNR, SSIM, LPIPS etc. Even efforts to make super-resolution techniques efficient are evaluated on these benchmarks, completely ignoring the fact that the final observers are human beings who have an extremely compressive visual system, and the practical goal is to deliver high quality to human observers. We design a robust and very fast perceptual framework that controls the quality of super-resolution in regards to the limitations of human vision, such that computational resources are greatly reduced without any noticeable degradation in visual quality. Our results demonstrate that even with FLOP reductions by factors of-and-greater than 3x, our perceptual algorithm minimally controls the super-resolution output such that the final result is not visually distinguishable from the output of a fully resourced deep network. We optimistically believe that our pipeline can serve as a promising direction for realizing real-time perceptually optimized and controlled image/video enhancement and generative synthesis.

### v2

@@ B.2 Video Content
B.3 Information about participants
All participants were recruited from CS department, aged 20 to 30. The network branching user study was conducted with a total of 15 participants (4F, 11M). The channel depth user study was conducted with a sample of 9 participants (3F, 6M)). Finally, the network branching applied to videos user study was conducted with 14 participants (3F, 11M). They had normal or corrected-to-normal vision and were unaware of the experiment’s purpose.
C FLOPS for evaluation and performance bottleneck
We decided to use the average FLOPS for efficiency due to its universality, independence from specific machine characteristics and implementation, as well as the fact that this measure has been used in similar studies [23, 19, 53, 48]. In our method patches that require the use of a network with larger size/capacity could be considered as a bottleneck, and in such a case, using the average FLOPS does not convey this information. A wall clock time elapsed could be considered a better measure of performance, for our method we would like to stress that the execution of the full network will not necessarily be the bottleneck. In scenarios where the number of processors is smaller than the number of patches, scheduling will take care of evenly distributing the load. A solution to further distribute the load more evenly would be to consider multiple consecutive frames for processing. Furthermore, it is possible that no patch in an image requires full network, and the full network will never be used.
D Framework input
The input patch size (size of patches into which the image is divided, as shown in Fig. 3) to our HVPF is equivalent to the size of the receptive field of the upsampling model employed. In the case of the VDSR network, the input patch is $\frac{40}{k}\times\frac{40}{k}$ pixels. This is due to the fact that, during the HVPF prediction, the low-resolution image is being considered. Before being conveyed to the VDSR network, the LR image is upsampled through bicubic interpolation. Consequently, in the event of $\times$4 upsampling, the input to the HVPF is 10x10 pixels, corresponding to a patch of 40x40 pixels in the image upsampled with bicubic interpolation. In the case of the EDSR network, the input patch is 48 $\times$ 48 pixels, taken from the low-resolution image. As no prior upsampling is involved in the input image, there is no need to consider a lower-size patch. In certain instances, utilizing input patches smaller than the neural network’s receptive field may be advantageous, particularly in the context of smaller images.
E Extension - AR/VR Display
Contrast sensitivity models such as the StelaCSF appropriately model human contrast perception as a function of eccentricity, and thus can extend our model to account for acuity across the visual field. Modern VR/AR headsets have built in eye-trackers that can be used to control our framework. In Fig. 9, we present some preliminary results for our quality map estimation with different gaze positions on the screen. The top row shows the eccentricity map relative to the gaze location, and the bottom row shows how our prediction for required SR quality varies. As anticipated, our perceptual model predicts that higher quality resolution will be used when the user is looking, while for areas in the periphery, our model predicts that the lowest up-sampling quality will be used. The main application is rendering in a lower resolution throughout the field of view, and then up-sampling the rendering for real-time VR/AR displays using our technique.
Figure 9: Our model predictions based on gaze position with $\times 4$ super-resolution. In the first column, we have the original image and the corresponding quality map. In the other columns we have on top the eccentricity map expressed in degrees and bottom we have the corresponding quality map.
F More Qualitative Results
This section presents additional qualitative results, demonstrating our perceptual model predictions. Fig. 10 shows the maps generated by our method, illustrating the selective deployment of higher-quality reconstruction networks in regions of greater detail and contrast, and the use of lower-quality reconstruction networks in areas with less detail and contrast. A comparison of the SR results of the proposed method with those of the original networks is presented in Fig. 11. Furthermore, visual SR results of the proposed method and of the original network is also presented in Fig. 12,13,14,15.
Figure 10: Our method quality map prediction results for $\times 4$ upsampling.
@@ F More Qualitative Results
(a) Original Network
(b) Our method
Figure 12: The SR result ($\times$4) of VDSR network and our method applied to VDSR network for image DIV2K-0885.
(a) Original Network
(b) Our method
Figure 13: The SR result ($\times$4) of VDSR network and our method applied to VDSR network for image DIV2K-0878.
(a) Original Network
(b) Our method
Figure 14: The SR result ($\times$4) of VDSR network and our method applied to VDSR network for image DIV2K-0850.
(a) Original Network
(b) Our method
Figure 15: The SR result ($\times$4) of VDSR network and our method applied to VDSR network for image DIV2K-0815.